OpenStudy (sadasdasdsadas):

What is the solution to the linear equation? 4b + 6 = 2 – b + 4

OpenStudy (kaithang):

b = 0

OpenStudy (kaithang):

Simplifying 4b + 6 = 2 + -1b + 4 Reorder the terms: 6 + 4b = 2 + -1b + 4 Reorder the terms: 6 + 4b = 2 + 4 + -1b Combine like terms: 2 + 4 = 6 6 + 4b = 6 + -1b Add '-6' to each side of the equation. 6 + -6 + 4b = 6 + -6 + -1b Combine like terms: 6 + -6 = 0 0 + 4b = 6 + -6 + -1b 4b = 6 + -6 + -1b Combine like terms: 6 + -6 = 0 4b = 0 + -1b 4b = -1b Solving 4b = -1b Solving for variable 'b'. Move all terms containing b to the left, all other terms to the right. Add 'b' to each side of the equation. 4b + b = -1b + b Combine like terms: 4b + b = 5b 5b = -1b + b Combine like terms: -1b + b = 0 5b = 0 Divide each side by '5'. b = 0 Simplifying b = 0